OpenStudy (anonymous):

Basic Algebra question:

OpenStudy (anonymous):

@Maria195 . 4(s+3) = 8

OpenStudy (anonymous):

Do you know how to distribute?

OpenStudy (anonymous):

I think so. Can i type the answer when I'm done and you'll me if i'm right?

OpenStudy (anonymous):

Yeah :)

OpenStudy (anonymous):

Simplifying 4(s + 3) = 8 Reorder the terms: 4(3 + s) = 8 (3 * 4 + s * 4) = 8 (12 + 4s) = 8 Solving 12 + 4s = 8 Solving for variable 's'. Move all terms containing s to the left, all other terms to the right. Add '-12' to each side of the equation. 12 + -12 + 4s = 8 + -12 Combine like terms: 12 + -12 = 0 0 + 4s = 8 + -12 4s = 8 + -12 Combine like terms: 8 + -12 = -4 4s = -4 Divide each side by '4'. s = -1 Simplifying s = -1

OpenStudy (anonymous):

Ugh I need help with this one. @Maria195

OpenStudy (anonymous):

Thanks @Okean

OpenStudy (anonymous):